40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14211     Entry time: Sun Sep 23 17:38:48 2018
Author: yuki 
Type: Update 
Category: ASC 
Subject: Alignment of AUX Y end green beam was recovered 

[ Yuki, Koji, Gautam ]

An alignment of AUX Y end green beam was bad. With Koji and Gautam's advice, it was recovered on Friday. The maximum value of TRY was about 0.5.

ELOG V3.1.3-